# Large-scale Urban Facility Location Selection with Knowledge-informed Reinforcement Learning

### v1

@@ 1. Introduction
In this paper, we present a deep reinforcement learning (DRL) method to solve the large-scale urban facility location problem, which is able to achieve comparable accessibility performance to commercial solvers while generating solutions with up to 1000 times speedup. Specifically, we construct a graph of urban with geospatial information, based on which we distill the essential swap operation in local search, replacing inefficient exploration with rapid edge selections which informed by a GNN tailored to the FLP. Additionally, we integrate tabu search (Rolland et al., 1997) and node benefits (Resende and Werneck, 2007) into the graph via dynamic wiring to avoid redundant and low-quality solutions. Finally, to address practical urban FLP with multiple types of facilities, we design a divide-and-conquer algorithm to solve the problem with two stages, which first accomplishes single facility solutions independently, then conducts adjustments, both using the unified SWAP operation and GNN model.
@@ 2. Problem Formulation
Given the total number of facilities $P$, candidate locations for facilities $\mathcal{N}=\{N_{1},N_{2},\cdots,N_{n}\}$, and corresponding demands =$\mathcal{H}=\{h_{1},h_{2},\cdots,h_{n}\}$, the FLP aims to maximize the accessibility of facilities. Different from the FLP, which focuses on a single type of facility, we investigate Multi-Facility Location Problem (MFLP), which involves $\mathcal{K}$ different types of facilities which are distinguished by different superscripts, and the type $k$ facility have to be placed in $P^{k}$ different regions among $\mathcal{N}$, with the corresponding residential demands are denoted as $\mathcal{H}^{k}=\{h_{1}^{k},\ldots,h_{n}^{k}\}$. The MFLP aims to optimize accessibility to all types of facilities, which can be measured by the access cost (AC) as follows:
@@
where $X_{i}^{p}$ is an indicator variable to whether a facility of type $p$ is placed at $N_{i}$.
@@ Acknowledgements.
This work is supported in part by National Key Research and Development Program of China under grant 2022ZD0116402 and National Natural Science Foundation of China under grant U22B2057 and 62272260. This work is also supported in part by Beijing National Research Center for Information Science and Technology (BNRist).

### v2

@@ 1. Introduction
In this paper, we present a deep reinforcement learning (DRL) method to solve the large-scale urban facility location problem, which is able to achieve comparable accessibility performance to commercial solvers while generating solutions with up to 1000 times speedup. Specifically, we construct an urban graph with geospatial information, based on which we distill the essential swap operation in local search, replacing inefficient exploration with rapid edge selections which are informed by a GNN tailored to the FLP. Additionally, we integrate tabu search (Rolland et al., 1997) and node benefits (Resende and Werneck, 2007) into the graph via dynamic wiring to avoid redundant and low-quality solutions. Finally, to address practical urban FLP with multiple types of facilities, we design a divide-and-conquer algorithm to solve the problem with two stages, which first accomplishes single facility solutions independently, then conducts adjustments, both using the unified SWAP operation and GNN model.
@@ 2. Problem Formulation
Given the total number of facilities $P$, candidate locations for facilities $\mathcal{N}=\{N_{1},N_{2},\cdots,N_{n}\}$, and corresponding demands =$\mathcal{H}=\{h_{1},h_{2},\cdots,h_{n}\}$, the FLP aims to maximize the accessibility of facilities. Different from the FLP, which focuses on a single type of facility, we investigate Multi-Facility Location Problem (MFLP), which involves $\mathcal{K}$ different types of facilities which are distinguished by different superscripts, and the type $k$ facility has to be placed in $P^{k}$ different regions among $\mathcal{N}$, with the corresponding residential demands denoted as $\mathcal{H}^{k}=\{h_{1}^{k},\ldots,h_{n}^{k}\}$. The MFLP aims to optimize accessibility to all types of facilities, which can be measured by the access cost (AC) as follows:
@@
where $X_{i}^{p}$ is an indicator variable for whether a facility of type $p$ is placed at $N_{i}$.
@@ Acknowledgements.
This work is supported in part by National Key Research and Development Program of China under grant 2022ZD0116402 and National Natural Science Foundation of China under grant U22B2057 and 62272260. This work is also supported in part by Tsinghua University-Toyota Research Center and Beijing National Research Center for Information Science and Technology (BNRist).